Teacher

William Astle Februrary 15, 2004

Notes

This work is copyright (c) 2004 by William Astle.

This work is licensed under the Creative Commons Attribution-NoDerivs-NonCommercial License. To view a copy of this license, visit http://creativecommons.org/licenses/by-nd-nc/1.0/ or send a letter to Creative Commons, 559 Nathan Abbott Way, Standford, California 94305, USA.

I am also greatly interested in any feedback anyone may have. You can send such feedback by email to lost@l-w.net. You may also be able to find more works by me at my web site at http://lost.l-w.net/ so please stop by.

Teacher

The scenery was breathtaking. In the years since the collapse, the trappings of a post modern society had all but vanished from the grasslands and plains of North America. Where there had once been long straight roads with automobiles screaming from place to place now exhibited a quietness. Not so much a desolation as a rebirth. What had once been square after square of crop land was now a rich collection of vegetation and wildlife. There was no question that the land would never return to precisely the same state it had been in before mass agriculture and dense human population, yet it was clearly recovering nicely.

Indeed, as Nor walked down the pathway from his cabin in a small wood, he realized that the wood itself had been encroaching on the grasslands over the time he had spent there. It was amazing how long slow changes just suddenly become apparent to an observer. Here he had been living for the past three centuries and he had not realized how much his little forest had expanded. How many times he had walked this very path in those centuries and never noticed the changes, the trees sprouting, growing, dying, in the perpetual cycle of life and death that constituted nature.

As he walked toward the village that still lay several kilometres from the wood, he came to a decision. He had closeted himself away from the world for too long. There had been sufficient time for humanity to recover from the collapse and now it was time to find out what they had done with the place. Feeling as though a weight had been lifted from his shoulders by this decision, he began whistling a tune to the time of his footsteps. Strange that he felt so liberated by this decision; after all, he had been happy in his cabin in the wood for so long. Perhaps a longing for change had been building within him similar to the way his wood had expanded. That must have been it, he decided.

It was shortly past midday when he arrived in the village. It was a remarkably orderly collection of buildings on either side of the pathway. Although he had to admit that the pathway had widened somewhat as he approached the village. He supposed it would qualify as a road here. He observed a few children frolicking around at the far end of the village as he entered from the west. Yet there were no adults in sight. Then he heard a cheer from a nearby building which bore a sign that read ”Tavern”. It was the only building which bore a sign of any kind. He wondered why that was.

As he entered the tavern, he realized that every adult who lived in the village was gathered in the room.. They nearly filled it to capacity but not so much that it was uncomfortable. He scanned the crowd looking for faces he recognized and spotted none. It hadn’t been that long since he’d been to the village, had it? As he thought about it he realized that it had indeed been a long time. Perhaps as much as a century. Strange that he hadn’t realized it until now. He realized at this point that he truly did need to get out into the world again; it was passing him by as he idly passed his days toying with his talents.

Seeing an unoccupied seat at a table near the door, he removed his pack and sat down, content to watch until he figured out what was happening. He had apparently arrived during a gap in the program as shortly a somewhat tall man with graying hair appeared on the slight stage opposite the entrance. Immediately the crowd silenced and the man began, ”In the years since the collapse, we have faced hardships that would have daunted the most indomitable pioneer. There isn’t a family among us that wasn’t devastated by the plague and later famine. Yet we survived, as did many like us.

”In the time leading up to the collapse, there were millions of people within a week’s walk from here. Everywhere there were roads and farms and towns and cities. There were great machines of burden lumbering noisily across the landscape. The natural world was slowly devastated by human activity until it finally cried out in agony. And we survived, we lucky few.

”When the collapse came, and less than one out of a thousand people survived, things looked hopeless. Then, just as the survivors came to terms with their loneliness, the land’s rebellion slammed them down again. And they survived, again reduced. Then, fighting between the tribes that came together threatened to finish the destruction that plague and famine failed to. And still, humanity survived.

”Now, we grow complacent. We are content to continue our lives exactly as we have since that group of survivors founded this village. We no longer strive for a better life. We no longer strive to learn. Yet it was the very knowledge they had gained that allowed our ancestors to survive the collapse and its aftermath! We MUST teach our children. We MUST encourage them to learn. And to ask questions. If we do not, we doom ourselves to be forgotten as time moves on without us! Yet if we leave it to the likes of Harvey, we would be left without even the ability to read! We MUST fight back. We must not let the likes of Harvey steal our future!”

As the old man stepped down, Nor observed several people nodding to themselves but nobody said anything. He wondered what had happened to these people. To the vibrant zest for life they had once demonstrated. He hadn’t long to wait for an answer to that question.

A young man, clean shaven and of average height, at least average for the villagers - Nor had been out of circulation long enough to be unsure of anything that would be described as average - took up residence on the stage. In contrast to the practical loose clothing of the old man, this youngster was dressed in what appeared to be a uniform of some kind, although no insignia of any sort were obvious. There was a definite change in the ambience of the room at this man’s appearance. Everyone was most certainly afraid of this man.

”Frank has made my point for me,” the young man began. ”He is right when he says that after the famine, war between tribes nearly wiped everyone out. And, mark my words war will come again. We must be ready for it at any cost. We must not let our village be destroyed by outsiders! We must train every man, woman, and child who is capable of fighting!

”While we have been comfortable here, in Norstown and Calgary and Red Deer, populations have been booming. How long before they start looking for more places to expand? The so called Free Alliance has been subsuming every town, village, nomadic tribe, and earthworm it came across. How long before they send in their forces to take over here? Do any of you wish to be told how to live your lives by bureaucrats five hundred kilometres distant?”

Nor had heard enough and unobtrusively ducked out. He had no desire to be painted as an evil conqueror should someone realize he was one of the man’s outsiders. He assumed that the young man was Harvey and he sincerely hoped that Harvey was simply trying to take over his own little piece of the world and that what he had implied about the Free Alliance was not true. Nor found himself a comfortable place to sit where he was unlikely to be noticed. Then he decided he should keep an eye on the proceedings in the tavern so he conjured a small globe that was barely visible and sent it into the tavern.

As he listened to the meeting via his magical spy, he pondered the state of the village. It was clear that Harvey had held some amount of sway for some time. Or perhaps something else had happened to the village to allow him to gain such power over the people. This, perhaps, explained the lack of signs over the other apparent business establishments. As he pondered the village’s predicament, he became aware that there was a commotion in the tavern. Apparently the old man had decided to argue some more and the scene was turning ugly. The small amount of support he’d had initially seemed to have completely eroded. Before Nor could decide what to do, the old man was dragged off into a back room to the cheers of the crowd. As much as Nor disagreed with the treatment of the old man, that very treatment had likely saved his life judging by the look of the crowd.

After a brief listen to the meeting to determine that the old man was likely doomed, he sprang to his feet and carefully, and swiftly, made his way around the back of the tavern. Once there, he found the back door was barred but there were no guards. Apparently the old man was not considered a threat. Then again, if everyone was in the meeting and village was in the middle of nowhere, who would there be to spring the fellow? Certain that nobody was waiting inside the door, he reached out with his magic and carefully lifted the bar and opened the door. Then he quietly entered the back corridor. Fortunately there was only one door which was locked and inside he found the old man. Fortunately the old man had his back turned to the door as Nor entered quietly. He crossed the room and placed his hand over the old man’s mouth and said, ”Please, not a sound.” When the old man nodded, Nor removed his hand.

The old man turned around and said, ”Who are you and what are you doing?”

”Never mind who I am. We have to get you out of here - that crowd out there is getting ugly.”

The old man hesitated for a moment then said, ”Lead the way.”

Nor led the old man out the back door, making sure to leave all the doors latched just as they had been. He suppressed a chuckle at the thought of the look on Harvey’s face when he found that his prisoner was missing as though by magic. Of course, that’s just what it had been.

Once they were clear of the tavern, Nor asked, ”Is there somewhere we can hide for a while?”

The old man thought for a moment then said, ”There’s a cellar in the old smithy. I don’t believe anyone would think to look there. Follow me.”

Several minutes later, the two of them had installed themselves in a surprisingly comfortable cellar. The old man clearly had questions for Nor but apparently had thought it would be better to save them for a more convenient time. This turned out to be wise as not long after they had hidden, the tavern emptied with people looking everywhere for the old man. As the old man had suggested, they didn’t think to look in the cellar.

After the search had died down, the old man said, ”Not that I’m not grateful, but who are you and why did you help me?”

Nor debated for a moment how much he should say then settled for saying, ”I’m a traveller who happened to stumble into that meeting back there right about when you made your speech about hardship. Then I heard the first bit of what that little warmonger had to say. I ducked out before anyone noticed a stranger in the meeting but I continued to listen in and when I heard the lynch mob forming, I decided to get you out of there.”

”I see. So what should I call you? I’m Frank by the way.”

”It’s been so long since I needed a name I’ve forgotten what mine is.” It wasn’t quite true but Nor felt it might lend some credence to the traveller story.

”Been away from people that long, eh? You must be older than you look.” ”You don’t know the half of it.” ”Okay, I’ll buy that. I’ll call you John if you don’t mind.” ”John. Hmm. Okay, I can live with that.” ”So where are you from? I can’t place your accent.” ”Around. I’ve been wandering for as long as I can remember.” Again, not

quite true. ”I’ve spent most of my time in the wilds to the north. It’s been some time since I anywhere near the Free Alliance or any other sizable population.”

”You said you heard my speech.”

”I did. It resonated. Too bad most folks here don’t seem too interested.” ”What’s your stake in all this anyway? I mean, why would a stranger take an

interest in the goings on in a small village on the edge of nowhere?” Something in Nor’s expression, or maybe his eyes, caused Frank to pause and take on a solemn expression. ”You have the look of a man who has seen too much, all of a sudden. I’ll not pry further into your reasons. You saved my life and that will suffice for now.”

”I thank you for your consideration.”

Frank and Nor remained in the cellar until nightfall.. They passed the time discussing random topics but not really discussing anything important. During that time, each man got a reasonable feel for the other and the bonds of friendship began to form. As night fell, the subject of what they would do came up.

”Well, it looks like my attempt to build up a following among this group has failed. I don’t expect it would be wise to stick around.” A note of sadness tinged Frank’s voice.

”I don’t know that it failed so much. You were getting through to a few of them until things went sidewise there. They haven’t felt the pain of oppression so they do not understand what it is they are doing by giving in to the fear. In time they will, and now that you’ve been run out of town, or, more likely, declared dead, some folks may think back to your speech as things go from bad to worse.”

”You sound like you’ve seen it before. And maybe you have - there’s something about you that I can’t put my finger on. And you make sense, but that still begs the question of what do we do now, for I’m certain you wouldn’t be welcome here any more than I am.”

”I might not be welcome, but I might be tolerated. However, that would require playing along and I don’t think I can swallow my principles just to live somewhere. I suppose we’ll end up wandering for a while.”

The two of them thought about that for a moment, then Frank said, ”If we’re going to wander, we’ll need some sort of cover. I wonder what would work?”

The discussion went on for a while. Then, still not having resolved the issue, they decided it was time to move out. In an unspoken agreement, they started out toward Norstown and the centre of the local piece of civilization. After a while, Frank commented, ”You seem to know your way quite well given the lack of a well defined path.” Nor made no reply and they continued in silence for the rest of the night.

Come dawn, they were some twenty kilometres from the village and, according to Frank, at least forty kilometres from the next village. They rested for a while and Nor shared some food out from his pack. Then they pressed on. Neither man felt comfortable being anywhere near the village. They continued in silence for the rest of the day, stopping around midday for another brief rest and food. By nightfall, they were about a day’s walk from the next village.

Nightfall the following day brought them to the village. This was a far different sight than Nor’s arrival at Frank’s village. Not only were there children playing in the streets but there were people going about their evening business. There were signs over most of the shops. While people looked at them with some surprise, and some suspicion, this was only natural in a world where travellers were rare. Mixed in with this was an air of excitement that spread like wildfire through the population of the village. Indeed, the village would perhaps qualify as a town; the collection of buildings reached back some distance on either side of the pathway they had picked up about half a day out. In fact, all along the way there were remnants of what was once a road between this town and Frank’s village. This one even had the remnants of a sign on the edge of the townsite which likely had once indicated a name for the town. The reason for its neglect was fairly obvious - there was nothing really beyond that edge of the town.

The two strangers were quickly herded into the local inn. Nor found it odd that even in a world where travellers were few and far between townsfolk built inns. Then again, in an agricultural setting where a trip to town was the better part of a day, the farmers needed somewhere to stay when they were in town. In the inn, Frank produced a small amount of what Nor assumed passed for currency these days and they had rooms for the night and food. After quickly getting settled in their rooms, they returned to the common room and spent the evening in the company of the other patrons. Most people seemed curious about strangers; they hadn’t seen many and those they had not stopped in the town. One conversation was particularly interesting to Nor, however.

It was getting somewhat late and Nor was just considering retiring to his room when a young boy, perhaps thirteen years of age, approached him. Well, not quite approach so much as slowly worked up the courage to take each step closer until he took a seat on the bench opposite Nor. Then he simply looked at Nor. After a while, Nor said, ”You’re up rather late aren’t you?”

Suddenly, the boy seemed to lose all fear. Perhaps the mysterious stranger had suddenly become a man rather than a supernatural being by speaking in such a normal manner. The boy said, ”My father said I could stay up late tonight on account of you and your friend arriving in town. What’s it like, travelling around? Where do you come from? What’s your name? Where are you going? Why are you travelling?”

Nor grinned. Then he replied, ”Tiring. A long way from here. John. Wherever I end up. That’s a long story.” Then he near burst out laughing as the boy tried to work out his answer. After a moment, it was almost as if a light bulb went on over the boy’s head as comprehension dawned in his eyes. Then Nor continued, ”What’s it like living here? What’s the name of this town? What’s your name? What are you going to be when you grow up. Why aren’t you travelling?”

The boy immediately responded, ”It’s great. Drysden. Frederick. I don’t know yet. I’m not old enough.”

Nor thought about the response for a split second then truly did laugh, a great belly laugh. Everyone in the common room stopped what they were doing and turned toward Nor and Frederick. After a moment, they all resumed what they were doing and Nor said, ”I like you, Frederick. You’ve got a sharp mind in that head of yours. You must do well in your schooling.”

”What is ’schooling’?”

Nor was not surprised by the response, but was somewhat disappointed. He responded, ”It is where you learn things about the world and other things.”

”Oh, like when father started showing me how to read and write?”

”Exactly like that, yes. Now, old John is tired from all the travelling to get here so why don’t you run along now?”

”Okay.” Just as he was about to disappear from view, Frederick turned waved at Nor. Then he turned back and disappeared into the back of the inn. Shortly thereafter, Nor retired to his room to ponder what the boy had said. A faint glimmering of an idea was forming in his mind and he needed to sleep on it before he discussed it with Frank.

In the morning, Frank and Nor were talking over breakfast. The inn was deserted since most folks were at home. ”I notice you were having a rather good time last night,” Frank said.

”I did. It’s been a long time since I spent time just talking to people. In fact, I had an idea last night that I’d like to bounce off you.”

”Okay, go ahead.”

”It occurred to me last night that the more people learn, and the more they are able to learn for themselves, the less likely people will fall for people like Harvey. Of course, you cannot force them to think for themselves, but you can at least give them the tools to do that thinking. You agree so far?”

”It makes sense, yes.”

”Now, most folks in these outlying areas do not have access to much in the way of learning. Books are rare and expensive and most schooling is done at home and limited by how much time and knowledge the parents have available. The rest of what people learn is from experience. You would agree with that, too?”

”I would, yes. I think I see where you’re going, but please continue.”

”It occurred to me last night that if someone were to take their time and teach people, it could go a long way toward bringing forth an enlightened society.”

”But what would you teach people? I mean, one person would have to spend a rather long time in each location to teach people anything particularly useful.”

”That’s the beauty of it. Instead of teaching people to read, write, think, and so on, which would be, as you observe, impractical. No, what I intend to do, instead, is to train teachers. And to do guest lectures when I’m in town. Perhaps if we can get an education system in place in these outlying areas, we can start a thought revolution.”

”What you propose could take decades. Are you prepared to dedicate your life to this idea of yours?”

”I’ll let you in on something. I’ve seen moving images of the world before the collapse. That world had wonders well beyond what we have now. It was possible to travel from your village to Norstown in a single day, or less. So much has been lost in the times that have followed. Yet so much was saved by those with foresight. There are great caches of books and other items waiting for a society that can understand them. Having seen the wonders of just one of those caches, I cannot live in a society that is progressing away from them. Yet these wonders had their own demons and I have seen the possible destruction they can cause. If the Harveys of the world gain control, it will not prevent the discovery of these caches, nor the redevelopment of those wonders, but in their hands, they will become terrors. Can you imagine the entire population of the earth dying in the aftermath of a single explosive device? There were chilling descriptions of just such an event from not long before the collapse. Knowing what I know, what kind of person would I be if I did not do everything in my power to prevent that sort of catastrophe? What sort of person would I be if I did not at least try to elevate our people above simply eking out a living?”

Again, Frank was left with a thoughtful look in his eyes as he said, ”You speak with the voice of true dedication. I would love to see the fruits of your labour but, alas, I am an old man. I will have to content myself with travelling along with you to see your small victories.”

”I would be delighted to have you along, something familiar in an unfamiliar world. And plus, I need a guide. Someone who has at least a vague idea what’s around these parts. And your belief in me helps too.”

”It’s a monumental task you’ve taken upon yourself. And there are no guarantees that you will survive to complete it. While I know I will never see you accomplish your goal, I don’t know if it will be accomplished in your lifetime either.”

”Well, the nature of this task is that even if I can give it a good solid start, there’s a better than even chance it will gain momentum and take on a life of its own. There’s a risk of that happening anyway. In creating a race of knowledgeable, thinking beings, I could well create just the opposite of what I want. But I must make the effort. The only thing that guarantees failure in a task is not to try at all. So really, whether I live to finish or not, a difference will be made.”

”With that kind of belief, you will succeed.”

The two of them had little trouble convincing the town council that education was a good idea. Over the following year, Nor sought out residents of Drysden who had knowledge and skills that were useful and taught them how to teach. Meanwhile, Frank was busy arranging for a teaching venue and the appropriate structure for the teaching. Thus the first school of Drysden was created. Nor also spent much time circulating among the population demonstrating the wonders of knowledge. Then, in the spring following their arrival in Drysden, Nor gave a final lecture - this time on history, the collapse, its aftermath, the time before the collapse, and the coming times. He told the story of arts and sciences and the wonders they achieved. He told of the horrors that had been conceived and perpetrated. He spoke of humanity and values, of tolerance and understanding. He spoke for most of the day, and not a single person moved from their seats, not for food, water, or to relieve themselves. And not just the children were there. Almost the entire population of the village had turned out for Nor’s farewell speech. Most of them had to stand in the hall, or outside it. Yet by some miracle, all who wanted to could hear what he had to say. He spoke with such conviction that his words moved even the most skeptical and the most cynical. In the end, most were sad to see him leave, but most began to understand his mission that day. Little did he know that that day would go down in the history books as the major turning point on the long road to recovery from the collapse.

Over the following years, Nor continued to tour the outlying towns that still had not been subsumed by the Free Alliance. As he did so, he learned of the Alliance and its practices and began to understand how Harvey had come to his conclusions. Yet he could detect a level of democracy and freedom behind the apparently expansionist policies of the Alliance. He had no doubt it had started out well enough intentioned but as with many things, it had attained a life of its own. He also learned that many had tried to topple the corrupt ruling structure of the Alliance and as many had failed. Some spectacularly and others quietly. Some had met with limited success only to discover that the evil of the Alliance was akin to the mythical hydra.

It was not even a decade later when it became apparent how much of an impact Nor was having. As he and Frank entered one fairly large town just beyond the border of the Free Alliance, a peddler on the street treated them to stories of a mysterious teacher roaming the Outlands (as the towns beyond the Alliance had come to be called). This teacher had no name but where he went, people were forever changed. In addition, he treated them to tales of how some towns had managed to beat back the Alliance’s advance and remain free. And these towns were all towns that Nor had visited previously. Upon hearing this, the two of them adjourned to a quiet place and Frank commented, ”I never dreamed that even a small amount of success would be visible before I passed on yet here we are, not even a decade into your task, and it is clear that you are making a difference. If only everyone could be treated with such knowledge in their lifetime.”

”I couldn’t have done it without you, old friend. Without your support in the beginning, none of this would have been possible. Indeed, were it not for that incident at your village - what is the name of that village, anyway? I never did find out - I would never have been in a position where the idea even occurred to me.”

”I appreciate the credit you give me. Yet somehow, I think you would have managed to come up with it anyway, and probably met with similar success. If I learned anything about you in the past years, it is that.” Then Frank’s face took on a truly serious expression and he said, ”Alas, this old body is getting too old for all this traipsing all over the Outlands. I’ll be retiring from this little operation after this town. As far as I can recall, this is the last large town in the Outlands, but who can tell where people have moved to since my information was gained.”

”We both knew this day would come. Where will you retire?”

”I was thinking I would go back to Drysden and, perhaps, write my memoirs. Somehow, I think they will be valuable at some point. I only ask of you one thing before I retire to Drysden. I ask that you finally tell me of your past and how you happened to be in Rockywood to save my life that day.”

Nor thought about the request for some time, during which Frank merely watched him patiently. Then he came to a decision. ”What I am about to tell you must be kept strictly confidential. It must not be present in your memoirs and you must not tell another living soul.”

When Frank made the promise, Nor began his tale, ”I was born on New Year’s Day in the year 1979 (old style) in a town that was then known as Olds. My name was Jeremy Martin Smith.” He held nothing back as he told Frank of his life before the collapse, of his eventful trip to the arctic, and his later discovery that he was not aging, and that he was not falling ill. He told of his founding of a cache of knowledge in the decades leading up to the collapse, and of his later survival of the epidemic, then the drought and famine, then the tribal wars. He told Frank of his choice of the name Nor and the founding of Norstown on the old townsite of Olds. He told of his later retirement to the woods just past Rockywood, also apparently known as Rockywood. He told of his initiation into the magical arts, and his isolated studies in his cabin in Rockywood. And of the wanderlust that had brought him down to the village on that fateful day. Frank listened in ever increasing awe and understanding - true understanding - slowly filled his eyes. There was no trace of disbelief.

When Nor finished telling his tale, Frank simply said, ”Thank you.”

Nor’s customary farewell speech when they finished their task in this town, which was known as Southern Hill, was the most moving of all his guest lectures, past and future, as he paid special tribute to his retiring friend. As with every speech to date, everyone who could come did. But this one left an even greater mark on those who heard it. Although most could not identify it, those who did were astounded by the depth of friendship their teacher was able to show for his partner and even more astounded at how clearly this feeling projected into the crowd. Frank, too, detected it and understood. His friend was bidding him farewell and telling him that they likely would never meet again. Yet neither of them had any regrets and they had said all they needed to prior to the speech. So the following day, without a word, they walked out of Southern Hill in opposite directions. Neither one looked back.

The speech followed by the wordless leave-taking the following day a truly life changing experience for all who witnessed it. It was nearly as life changing for those who heard it second hand or third hand. Eventually, word of it began to filter into the towns inside the Free Alliance to compound with the news of the teacher’s continued exploits. For the second time in less than a decade, the course of history was turned although the implications of this change would be far more profound than anyone would have predicted.